# Efficient Computation of Voronoi Diagrams Using Point-in-Cell Tests

### v1

@@ 4.4.2. Propagation
Figure 6. A modified solution for generating intersection $\mathcal{P}$ of a triangle/tetrahedron $\mathcal{T}$ and the Voronoi cell $\Omega_{i}$. After $\mathcal{P}$ is initialized as $\mathcal{T}$, it is clipped iteratively. In each iteration, we select a corner point of $\mathcal{P}$ that is outside $\Omega_{i}$ and build a virtual edge (pink dotted line) by connecting $\mathbf{v}_{i}$ and the corner point, based on which we can find a bisector by using the proposed edge-based search scheme to clip $\mathcal{P}$. After a few iterations, the final $\mathcal{P}$ has no corner point being outside $\Omega_{i}$. The generation is independent for each site and suitable for parallel computing. $i=4$ in this figure.
@@
The last obstacle is the generation of $\mathcal{P}_{a,j}$, where $j$ is the index of the site popped from the queue in each iteration. If there is at least one corner point of $\mathcal{T}_{a}$ being in the Voronoi cell $\Omega_{j}$, Algorithm 1 can be directly applied to compute $\mathcal{P}_{a,j}$. Otherwise, no edge can be selected from $\mathcal{T}_{a}$ to launch clippings. Therefore, an intermediate polygon/polytope $\mathcal{G}_{a,j}$ that has at least one corner point being in $\Omega_{j}$ should be constructed before computing $\mathcal{P}_{a,j}$.
@@
5. Parallel Computation on GPUs
In this section, we extend our method to GPUs for the parallel computation of clipped or restricted Voronoi cells. To achieve this, an independent solution for generating each triangle/tetrahedron-cell intersection is required. The solution presented in 4.3 becomes inapplicable here. We therefore propose a modified version, which will serve as the foundation for the parallel algorithms introduced subsequently.
5.1. Modified generation of triangle/tetrahedron-cell intersection
Because each site must be in its own Voronoi cell, to generate the intersection $\mathcal{P}$ of a triangle/tetrahedron $\mathcal{T}$ and a Voronoi cell $\Omega_{i}$, we can build virtual edges by connecting the corresponding site $\mathbf{v}_{i}$ and the corner points of $\mathcal{T}$. These virtual edges can then be fed into the proposed edge-based search scheme to find bisectors. Specifically, $\mathcal{P}$ is first initialized as $\mathcal{T}$, and it is clipped iteratively to be smaller and smaller. In each iteration, two steps are executed: (1) Select a corner point $\mathbf{p}$ of $\mathcal{P}$, requiring that $\mathbf{p}\notin\Omega_{i}$; (2) Find a bisector of $\mathbf{v}_{i}$ and other site by using the edge-based search scheme on virtual edge $\mathbf{v}_{i}\mathbf{p}$, and use it to clip $\mathcal{P}$, the result is still denoted as $\mathcal{P}$. After a few iterations, the final $\mathcal{P}$ has no corner point being outside $\Omega_{i}$. Fig. 6 shows an illustration, and Algorithm 3 lists the pseudo-code. This modified generation is independent of each site and suitable for parallel computing.
0: a triangle/tetrahedron $\mathcal{T}$ and the site index ii.
@@ 5.1. Modified generation of triangle/tetrahedron-cell intersection
3:  find a bisector $B_{i,j}$ on virtual edge $\mathbf{v}_{i}\mathbf{p}$ (section 4.2);
@@
ALGORITHM 3 Modified generation of triangle/tetrahedron-cell intersection
It is worth pointing out that applying the modified generation of polygon/polytope-cell intersection in serial computing is not suggested, since it will result in efficiency loss. For generating the intersection $\mathcal{P}$ of $\mathcal{T}$ and $\Omega_{i}$, the original solution presented in 4.3 ensures that at least one of the new corner points introduced by each clipping must be in $\Omega_{i}$, which is not supported in the modified solution. Consequently, the number of the point-in-cell tests in the modified solution is more than that in the original solution.
5.2. Parallel computation on regular domains
If the given domain is regular with only a few corner points (e.g., a rectangle in $\mathbb{R}^{2}$ or a cube in $\mathbb{R}^{3}$), the parallel computation of clipped or restricted Voronoi cells is straightforward. We assign one thread per site, the main task of each thread is to iteratively clip the domain by using Algorithm 3.
As mentioned in section 2, the parallelized kNN-based method [Ray et al., 2018] may yield approximate results. In contrast, our method guarantees exact results because Algorithm 3 employs an iterative process that continues until all corner points are in the target Voronoi cell. Furthermore, similar to the kNN-based method, our parallel algorithm requires no inter-thread data synchronization on GPUs, enabling high parallelism in computation.
5.3. Parallel computation on irregular domains
If the given domain is irregular and contains many small triangles/tetrahedra, the parallel computation of clipped or restricted Voronoi diagrams becomes much more complicated. Here, we focus on the surface and 3D cases.
5.3.1. Surface case
@@
Specifically, for each triangle $\mathcal{T}_{a}\in\mathcal{M}$, we identify the nearest site $\mathbf{v}_{i}$ of its centroid, forming a pair $(\mathcal{T}_{a},\mathbf{v}_{i})$. These pairs can then be assigned to GPU threads for computing corresponding triangle-cell intersections. Based on positive edge flags of the intersections, additional triangle-cell pairs can be discovered, similar to the propagation scheme of Han et al. [Han et al., 2017]. The computation terminates when no more triangle-cell pairs are obtained. Algorithm 4 outlines the steps.
0: 3D sites $\mathcal{V}=\left\{\mathbf{v}_{i}\right\}_{i=1}^{n}$, a triangular mesh $\mathcal{M}=\left\{\mathcal{T}_{a}\right\}_{a=1}^{m}$.
@@
1: build a structure $\mathcal{S}$ of $\mathcal{V}$ for nearest neighbor search (section 6.2);
@@
5:  insert $(\mathcal{T}_{a},\mathbf{v}_{i})$ to $\mathcal{L}_{1}$ in atomic and label $(\mathcal{T}_{a},\mathbf{v}_{i})$;
@@
10:   compute the intersection $\mathcal{P}_{a,i}$ of $\mathcal{T}_{a}$ and $\Omega_{i}$ using Algorithm 3;
@@
13:    if $f$ has a positive flag $j$ and $(\mathcal{T}_{a},\mathbf{v}_{j})$ is unlabeled then
14:     insert $(\mathcal{T}_{a},\mathbf{v}_{j})$ to $\mathcal{L}_{2}$ in atomic and label $(\mathcal{T}_{a},\mathbf{v}_{j})$;
@@
ALGORITHM 4 Parallel computation of restricted Voronoi diagram
5.3.2. 3D case
The domain is composed of $m$ tetrahedra in this case. Because we expect exact results, the propagation scheme described in 5.3.1 is also applied to compute 3D clipped Voronoi cells. In fact, only a subset of the sites needs to be processed in this manner.
By computing restricted Voronoi cells, the given sites $\mathcal{V}$ can be divided into two categories [Yan et al., 2013]: (1) boundary sites $\mathcal{V}_{\mathcal{B}}$ with non-empty restricted Voronoi cells, and (2) interior sites $\mathcal{V}_{\mathcal{I}}=\mathcal{V}\setminus\mathcal{V}_{\mathcal{B}}$. For an interior site $\mathbf{v}_{i}\in\mathcal{V}_{\mathcal{I}}$, its domain-cell intersection is identical to the Voronoi cell itself, i.e., $\Omega_{i|\mathcal{M}}=\Omega_{i}$. In contrast, for a boundary site $\mathbf{v}_{j}\in\mathcal{V}_{\mathcal{B}}$, $\Omega_{j|\mathcal{M}}\subset\Omega_{j}$ holds. Thus, the parallel algorithm in 5.2 can be directly applied to interior sites.
The clipped Voronoi cells computation of boundary sites are as follows. For each boundary site $\mathbf{v}_{j}\in\mathcal{V}_{\mathcal{B}}$, its restricted Voronoi cell consists of multiple polygons, each originating from a boundary triangle $\Delta$ of the domain. Thus, a pair $(\mathcal{T}_{a},\mathbf{v}_{j})$ is formed, where $\mathcal{T}_{a}$ represents the incident tetrahedron of $\Delta$. After initializing a set of tetrahedron-cell pairs, the remaining steps follow those described in 5.3.1, but negative face flags of the generated tetrahedron-cell intersections are used to propagate tetrahedron-cell pairs. Algorithm 5 provides the detailed procedure.
0: 3D sites $\mathcal{V}=\left\{\mathbf{v}_{i}\right\}_{i=1}^{n}$, a tetrahedralized mesh $\mathcal{M}=\left\{\mathcal{T}_{a}\right\}_{a=1}^{m}$.
@@
1: build a structure $\mathcal{S}$ of $\mathcal{V}$ for nearest neighbor search (section 6.2);
// steps 2 $\sim$ 5 are for dividing sites
@@
3: compute restricted Voronoi cells $\left\{\mathcal{C}_{i}\right\}_{i=1}^{n}$ on $\mathcal{M}_{\mathcal{B}}$ using Algorithm 4;
4: boundary sites $\mathcal{V}_{\mathcal{B}}\leftarrow$ sites with non-empty restricted Voronoi cell;
5: interior sites $\mathcal{V}_{\mathcal{I}}\leftarrow\mathcal{V}\setminus\mathcal{V}_{\mathcal{B}}$;
6: given $\mathcal{M}$’s bounding box, compute cells of interior sites (section 5.2);
// the rest steps are for computing cells of boundary sites
@@
9:  for each polygon $\mathcal{A}\in\mathcal{C}_{j}$ do
10:   $\Delta\leftarrow$ the boundary triangle that $\mathcal{A}$ originates from;
@@
12:   insert $(\mathcal{T}_{a},\mathbf{v}_{j})$ to $\mathcal{L}_{1}$ in atomic, and label $(\mathcal{T}_{a},\mathbf{v}_{j})$;
@@
18:   generate the intersection $\mathcal{P}_{a,j}$ of $\mathcal{T}_{a}$ and $\Omega_{j}$ using Algorithm 3;
@@
21:    if $f$ has a negative flag (b-bthen
22:     $\mathcal{T}_{b}\leftarrow$ the neighboring tetrahedron of $\mathcal{T}_{a}$;
23:     if $(\mathcal{T}_{b},\mathbf{v}_{j})$ is unlabeled then
24:      insert $(\mathcal{T}_{b},\mathbf{v}_{j})$ to $\mathcal{L}_{2}$ in atomic, and label $(\mathcal{T}_{b},\mathbf{v}_{j})$;
25:     end if
26:    end if
27:   end for
28:  end for
29:  $\mathcal{L}_{1}\leftarrow\mathcal{L}_{2}$;
30:end while
ALGORITHM 5 Parallel computation of 3D clipped Voronoi diagram
6. Implementation Details
6.1. Improved edge-based bisector search
Given an edge $\mathbf{p}_{s}\mathbf{p}_{t},\mathbf{p}_{s}\in\Omega_{i}$, the original version of edge-based bisector search presented in 4.2 initiates the search from the end point $\mathbf{p}_{t}$. A larger number of Voronoi cells intersected with the edge, more point-in-cell tests are required. We aim to reduce the number of tests to further enhance the efficiency of our method.
@@ 6.1. Improved edge-based bisector search
Let $\mathcal{N}_{i}$ be the nearest neighbor of $\mathbf{v}_{i}$. The distance $\delta=\left\|\mathcal{N}_{i}-\mathbf{v}_{i}\right\|/2$ implies an approximate size of $\Omega_{i}$, which can be used to guide the search of point $\mathbf{g}$. Specifically, initiate $\lambda=2,\mathbf{g}=\mathbf{p}_{s}+\lambda\delta\frac{\mathbf{p}_{t}-\mathbf{p}_{s}}{\left\|\mathbf{p}_{t}-\mathbf{p}_{s}\right\|}$, if the nearest site of $\mathbf{g}$ is $\mathbf{v}_{i}$, we update $\lambda\leftarrow\lambda+1,\mathbf{g}$ and try again, otherwise, the search of $\mathbf{g}$ finishes.
@@
6.2. Nearest neighbor search
Our method heavily relies on the point-in-cell tests, each of which requires a nearest neighbor search. The KD-tree is a well-known data structure that provides efficient nearest neighbor search, which has been widely applied in many fields. However, its search efficiency remains insufficient for our method due to the massive number of required nearest neighbor searches. For example, generating the result in Fig. 1(left) required over 20.54 million searches. Therefore, improving search efficiency is crucial to our method.

### v2

@@ IV-D Computation of clipped or restricted Voronoi diagram
Figure 7: A modified solution for computing intersection $\mathcal{P}$ of a simplex $\mathcal{T}$ and a Voronoi cell $\Omega_{i}$. After $\mathcal{P}$ is initialized as $\mathcal{T}$, it is clipped iteratively. In each iteration, we select a corner point of $\mathcal{P}$ that is outside $\Omega_{i}$ and build a virtual edge (pink dotted line) by connecting $\mathbf{v}_{i}$ and the corner point, based on which we can find a bisector by using the proposed edge-based search scheme to clip $\mathcal{P}$. After a few iterations, the final $\mathcal{P}$ has no corner point being outside $\Omega_{i}$. The computation is independent for each site and suitable for parallel computing. $i=4$ in this figure.
IV-D1 First polygon/polytope
IV-D2 Propagation
It has to further compute the intersections of $\mathcal{T}_{a}$ and other Voronoi cells, the simplex-cell pairs can be obtained by using the edge/face flags of $\mathcal{P}_{a,i}$. Specifically, we set up a queue $Q$ to store the unvisited sites whose Voronoi cells are intersected with $\mathcal{T}_{a}$. After the first polygon/polytope $\mathcal{P}_{a,i}$ is generated, we traverse its all edges/faces. For each edge/face, if it has a positive flag $k$, meaning that $\Omega_{k}$ must intersect with $\mathcal{T}_{a}$, we add $\mathbf{v}_{k}$ to the queue. Iteratively, a site $\mathbf{v}_{j}$ is popped from $Q$, and we compute the intersection $\mathcal{P}_{a,j}$ of $\mathcal{T}_{a}$ and $\Omega_{j}$, which is then added to $\Omega_{j|\mathcal{M}}$. The queue $Q$ can be further updated by adding more unvisited sites using the edge/face flags of $\mathcal{P}_{a,j}$. The procedure is terminated once the queue is empty, and the decomposition on $\mathcal{T}_{a}$ is completed.
@@ IV-D2 Propagation
The last obstacle is the computation of $\mathcal{P}_{a,j}$, where $j$ is the index of the site popped from the queue in each iteration. If there is at least one corner point of $\mathcal{T}_{a}$ being in the Voronoi cell $\Omega_{j}$, Algorithm 1 can be directly applied to compute $\mathcal{P}_{a,j}$. Otherwise, no edge can be selected from $\mathcal{T}_{a}$ to launch clippings. Therefore, an intermediate polygon/polytope $\mathcal{G}_{a,j}$ that has at least one corner point being in $\Omega_{j}$ should be constructed before computing $\mathcal{P}_{a,j}$.
@@
Algorithm 3 Modified computation of simplex-cell intersection
0: a simplex $\mathcal{T}$ and the site index ii.
@@
3:  find a bisector $B_{i,j}$ on virtual edge $\mathbf{v}_{i}\mathbf{p}$ (Section IV-B);
@@
V Parallel Computation on GPUs
In this section, we extend our method to GPUs for the parallel computation of clipped or restricted Voronoi cells. To achieve this, an independent solution for generating each simplex-cell intersection is required. The solution presented in Section IV-C becomes inapplicable here. We therefore propose a modified version, which will serve as the foundation for the parallel algorithms introduced subsequently.
V-A Modified computation of simplex-cell intersection
Because each site must be in its own Voronoi cell, to generate the intersection $\mathcal{P}$ of a simplex $\mathcal{T}$ and a Voronoi cell $\Omega_{i}$, we can build virtual edges by connecting the corresponding site $\mathbf{v}_{i}$ and the corner points of $\mathcal{T}$. These virtual edges can then be fed into the proposed edge-based search scheme to find bisectors. Specifically, $\mathcal{P}$ is first initialized as $\mathcal{T}$, and it is clipped iteratively to be smaller and smaller. In each iteration, two steps are executed: (1) select a corner point $\mathbf{p}$ of $\mathcal{P}$, requiring that $\mathbf{p}\notin\Omega_{i}$; (2) find a bisector of $\mathbf{v}_{i}$ and other site by using the edge-based search scheme on virtual edge $\mathbf{v}_{i}\mathbf{p}$, and use it to clip $\mathcal{P}$, the result is still denoted as $\mathcal{P}$. After a few iterations, the final $\mathcal{P}$ has no corner point being outside $\Omega_{i}$. Fig. 7 shows an illustration, and Algorithm 3 lists the pseudo-code. This modified algorithm is independent of each site and suitable for parallel computing.
It is worth pointing out that applying the modified computation of simplex-cell intersection in serial computing is not suggested, since it will result in efficiency loss. For generating the intersection $\mathcal{P}$ of $\mathcal{T}$ and $\Omega_{i}$, the original solution presented in IV-C ensures that at least one of the new corner points introduced by each clipping must be in $\Omega_{i}$, which is not supported in the modified solution. Consequently, the number of the point-in-cell tests in the modified solution is more than that in the original solution.
V-B Parallel computation on simple domains
If the given domain is simple with only a few corner points (e.g., a cube in $\mathbb{R}^{3}$), the parallel computation of clipped or restricted Voronoi cells is straightforward. We assign one thread for each site, the main task of each thread is to iteratively clip the domain by using Algorithm 3.
Our parallel algorithm guarantees exact results because Algorithm 3 employs an iterative process that continues until all corner points are in the target Voronoi cell. In addition, similar to the kNN-based method, our parallel algorithm requires no inter-thread data synchronization on GPUs, enabling high parallelism in computation.
V-C Parallel computation on complex domains
If the given domain is complex and consists of a set of small simplices, the parallel computation of clipped or restricted Voronoi diagrams becomes much more complicated. Here, we focus on the surface and 3D cases.
V-C1 Surface case
@@
Specifically, for each triangle $\mathcal{T}_{a}\in\mathcal{M}$, we identify the nearest site $\mathbf{v}_{i}$ of its centroid, forming a pair $(\mathcal{T}_{a},\mathbf{v}_{i})$. These pairs can then be assigned to GPU threads for computing corresponding triangle-cell intersections. Based on positive edge flags of the intersections, additional triangle-cell pairs can be discovered, similar to the propagation scheme of Han et al. [26]. The computation terminates when no more triangle-cell pairs are obtained. Algorithm 4 outlines the steps.
Algorithm 4 Parallel computation of restricted VD
0: sites $\mathcal{V}=\left\{\mathbf{v}_{i}\right\}_{i=1}^{n}$, domain triangles $\mathcal{M}=\left\{\mathcal{T}_{a}\right\}_{a=1}^{m}$.
@@
1: build a structure $\mathcal{S}$ of $\mathcal{V}$ for nearest neighbor searches (section VI-B);
@@
5:  insert $(\mathcal{T}_{a},\mathbf{v}_{i})$ to $\mathcal{L}_{1}$ and label it in atomic;
@@
10:   compute intersection $\mathcal{P}_{a,i}$ of $\mathcal{T}_{a}$ and $\Omega_{i}$ (Alg. 3);
@@
13:    if $f$’s flag $j>0$ and $(\mathcal{T}_{a},\mathbf{v}_{j})$ is unlabeled then
14:     insert $(\mathcal{T}_{a},\mathbf{v}_{j})$ to $\mathcal{L}_{2}$ and label it in atomic;
@@
V-C2 3D case
[…]
By computing restricted Voronoi cells, the given sites $\mathcal{V}$ can be divided into two categories [47]: (1) boundary sites $\mathcal{V}_{\mathcal{B}}$ with non-empty restricted Voronoi cells, and (2) interior sites $\mathcal{V}_{\mathcal{I}}=\mathcal{V}\setminus\mathcal{V}_{\mathcal{B}}$. For an interior site $\mathbf{v}_{i}\in\mathcal{V}_{\mathcal{I}}$, its domain-cell intersection is identical to the Voronoi cell itself, i.e., $\Omega_{i|\mathcal{M}}=\Omega_{i}$. Thus, the parallel algorithm in Section V-B can be directly applied to interior sites.
For boundary sites, the parallel computation of clipped Voronoi cells are as follows. For each site $\mathbf{v}_{j}\in\mathcal{V}_{\mathcal{B}}$, its restricted Voronoi cell consists of multiple polygons, each originating from a boundary triangle $\Delta$ of the domain. Thus, a pair $(\mathcal{T}_{a},\mathbf{v}_{j})$ is formed, where $\mathcal{T}_{a}$ represents the incident tetrahedron of $\Delta$. After initializing a set of tetrahedron-cell pairs, the remaining steps follow those described in Section V-C1, but negative face flags of the generated tetrahedron-cell intersections are used to propagate tetrahedron-cell pairs. Algorithm 5 provides the detailed procedure.
Algorithm 5 Parallel computation of 3D clipped VD
0: sites $\mathcal{V}=\left\{\mathbf{v}_{i}\right\}_{i=1}^{n}$, domain tetrahedra $\mathcal{M}=\left\{\mathcal{T}_{a}\right\}_{a=1}^{m}$.
@@
1: build a structure $\mathcal{S}$ of $\mathcal{V}$ for nearest neighbor searches (section VI-B);// steps 2 $\sim$ 5 are for dividing sites
@@
3: compute restricted Voronoi cells $\left\{\mathcal{C}_{i}\right\}_{i=1}^{n}$ on $\mathcal{M}_{\mathcal{B}}$ (Alg. 4);
4: boundary sites $\mathcal{V}_{\mathcal{B}}\leftarrow\left\{\mathbf{v}_{i}\lvert\mathcal{C}_{i}\neq\emptyset\right\}$;
5: interior sites $\mathcal{V}_{\mathcal{I}}\leftarrow\mathcal{V}\setminus\mathcal{V}_{\mathcal{B}}$;// cells of interior sites
6: given $\mathcal{M}$’s bounding box, compute cells of interior sites (section V-B);// cells of boundary sites
@@
9:  for each polygon $\mathcal{H}\in\mathcal{C}_{j}$ do
10:   $\Delta\leftarrow$ the boundary triangle that $\mathcal{H}$ originates from;
@@
12:   insert $(\mathcal{T}_{a},\mathbf{v}_{j})$ to $\mathcal{L}_{1}$ and label it in atomic;
@@
18:   compute intersection $\mathcal{P}_{a,j}$ of $\mathcal{T}_{a}$ and $\Omega_{j}$ (Alg. 3);
@@
21:    if $f$’s flag (b)<0(-b)<0 and $(\mathcal{T}_{b},\mathbf{v}_{j})$ is unlabeled then
22:     insert $(\mathcal{T}_{b},\mathbf{v}_{j})$ to $\mathcal{L}_{2}$ and label it in atomic;
23:    end if
24:   end for
25:  end for
26:  $\mathcal{L}_{1}\leftarrow\mathcal{L}_{2}$;
27:end while
VI Implementation Details
VI-A Improved edge-based bisector search
Given an edge $\mathbf{p}_{s}\mathbf{p}_{t},\mathbf{p}_{s}\in\Omega_{i}$, the original version of edge-based bisector search presented in IV-B initiates the search from the end point $\mathbf{p}_{t}$. A larger number of Voronoi cells intersected with the edge, more point-in-cell tests are required. We aim to reduce the number of tests to further enhance the efficiency of our method.
@@ VI-A Improved edge-based bisector search
Let $\mathcal{N}_{i}$ be the nearest neighbor of $\mathbf{v}_{i}$ among all the sites. The distance $\delta=\left\|\mathcal{N}_{i}-\mathbf{v}_{i}\right\|/2$ implies an approximate size of $\Omega_{i}$, which can be used to guide the search of point $\mathbf{g}$. Specifically, initiate $\lambda=2,\mathbf{g}=\mathbf{p}_{s}+\lambda\delta\frac{\mathbf{p}_{t}-\mathbf{p}_{s}}{\left\|\mathbf{p}_{t}-\mathbf{p}_{s}\right\|}$, if the nearest site of $\mathbf{g}$ is $\mathbf{v}_{i}$, we increment $\lambda$ by 1 and repeat the search with the updated $\mathbf{g}$; otherwise, the search terminates.
@@
VI-B Nearest neighbor search
Our method heavily relies on the point-in-cell tests, each of which requires a nearest neighbor search. The KD-tree is a well-known data structure that provides efficient nearest neighbor search, which has been widely applied in many fields. However, its search efficiency remains insufficient for our method due to the massive number of required nearest neighbor searches. For example, generating 1 million Voronoi cells bounded by a cube required over 20.54 millions searches, where the sites are randomly sampled. Therefore, improving search efficiency is crucial to our method.